OpenStudy (anonymous):

Out of 1,100 discs tested,13 are defective.EStimate the number of defective discs in a batch of 41,000

OpenStudy (anonymous):

well i would divide 41,000 by 1,100 then multiply that by 13

OpenStudy (anonymous):

(13/1100) = (x/41000) 533000 = 1100x divide both sides by 1100